# Experimental Analysis of Efficiency of the Messaging Layer Security for Multiple Delivery Services

### v1

@@ 8 Conclusion and Future Work
In this work we have analysed the performance of MLS in an empirical setting. We analyse the protocol under different parameters such as Delivery Service, chosen paradigm or state of the Ratchet Tree. Our results show that theoretical claims of logarithmic complexity in communication cost do not necessarily manifest in practice.
Our work also demonstrates the relevance of practical considerations in the performance of MLS, as opposed to theoretical analysis. We have found that the most significant time cost - by multiple orders of magnitude - is message distribution through the Delivery Service. Furthermore, the messages that introduce higher workload are Welcome and GroupInfo messages, as they contain the full Ratchet Tree. To the extent of our knowledge, no other work has attempted to increase the efficiency of these operations.
We also publish our execution environment as open source, which includes a configurable simulated MLS client and two different Delivery Services. We hope that it can help other researchers to perform their own simulations and that it can serve as a baseline for more complex simulation environments.
We plan on continuing to study the applicability of concepts defined specifically in MLS - such as the propose-and-commit paradigm or External Joins [13] - to general definitions of CGKA and their impact on the security properties and efficiency of these schemes.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgement
D.S. acknowledges support from Xunta de Galicia and the European Union (European Social Fund - ESF) scholarship [ED481A-2023-219].
This work is part of the project TED2021-130369B-C31 and TED2021-130492B-C21 funded by MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”.

### v2

@@ 8 Conclusion and Future Work
In this work we have analysed the performance of MLS in an empirical setting. We analyse the protocol under different parameters such as Delivery Service, chosen paradigm or state of the Ratchet Tree. Our results show that theoretical claims of logarithmic complexity in communication cost do not manifest in practice. Our work also demonstrates the relevance of practical considerations in the performance of MLS. We have found that the most significant time costs are message distribution through the Delivery Service and the storage of the Ratchet tree, which are not taken into account in theoretical analysis.
We also publish our execution environment as open-source, which includes a configurable emulated MLS client and two different Delivery Services. Our testbed can be employed by other researchers to perform in-depth studies of the behaviour of MLS in different environments, and can serve as a baseline for the development of more complex experiments. Our implementation can be easily adapted to execute other CGKA variants, such that their performance can be systematically compared to that of MLS. Future work should focus on implementing a more complex Delivery Service that can accurately represent real-world inconveniences and optimising the emulated client’s memory usage and parallelism to allow for larger experiments.
Acknowledgements
The work is funded by the Plan Complementario de Comunicaciones Cuánticas, Spanish Ministry of Science and Innovation(MICINN), Plan de Recuperación NextGeneration, European Union (PRTR-C17.I1, CITIC Ref.305.2022), and Regional Government of Galicia (Agencia Gallega de Innovación, GAIN, CITIC Ref. 306.2022). D.S. acknowledges support from Xunta de Galicia and the European Union (European Social Fund - ESF) scholarship [ED481A-2023-219]. We also acknowledge support from the Xunta de Galicia and the European Union (FEDER Galicia 2021-2027 Program) Ref. ED431B 2024/21, CITIC ED431G 2023/0.